Solve each absolute value equation:
4-3|x - 2|= -8

Answer by Edwin McCravy(20055)


Solve each absolute value equation: 
4-3|x - 2|= -8

Isolate the absolute value on the 
left:

Subtract 4 from both sides:

 -3(x - 2| = -8-4

 -3|x - 2| = -12

Divide both sides by -3

   |x - 2| = (-12)/(-3)

   |x - 2| = 4

To solve equations of the form

   |expression| = A

1. If A is negative there is no solution.
2. If A is 0, solve the equation expression = 0
3. if A is positive, solve both equations
 expression = A and expression = -A
 and get two solutions

Yours is case 3:

   |x - 2| = 4

so we solve both equations:

x - 2 = 4      x - 2 = -4
    x = 6          x = -2

Those are the two solutions.

Edwin